# Exact non-stationary solutions of the Euler equations in two and three dimensions

### v1

@@ 4.2.2. Curl eigenfields in a torus bundle
[…]
Here $\lambda=2m/\alpha$ in Theorem 2.2.
Proof.
The main thing is to compute curl eigenfields, which we know exist by Lemma 2.1. From the explicit form of the metric, we know that an orthogonal basis of $1$-forms is given by [^4] Here we use a non-orthonormal basis of $1$-forms, unlike in Theorem 4.6, since the $d\sigma^{k}$ are much simpler with this choice.
(4.18) $\sigma^{1}=dr,\qquad\sigma^{2}=d\theta+\frac{c}{\varphi(r)^{2}}\,dz,\qquad\sigma^{3}=dz,$
satisfying
$\lvert\sigma^{1}\rvert=1,\qquad\lvert\sigma^{2}\rvert=\frac{1}{\varphi(r)},\qquad\lvert\sigma^{3}\rvert=\frac{1}{\varphi^{\prime}(r)},$
and from these equations we can compute the Hodge star:
$\star\sigma^{1}=\frac{\lvert\sigma^{1}\rvert}{\lvert\sigma^{2}\rvert\lvert\sigma^{3}\rvert}\,\sigma^{2}\wedge\sigma^{3}=\varphi(r)\varphi^{\prime}(r)\,\sigma^{2}\wedge\sigma^{3},$
and similarly
$\star\sigma^{2}=\frac{\varphi^{\prime}(r)}{\varphi(r)}\,\sigma^{3}\wedge\sigma^{1},\qquad\star\sigma^{3}=\frac{\varphi(r)}{\varphi^{\prime}(r)}\,\sigma^{1}\wedge\sigma^{2}.$
We also have
$d\sigma^{1}=0,\qquad d\sigma^{2}=\frac{2c\varphi^{\prime}(r)}{\varphi(r)^{3}}\,\sigma^{3}\wedge\sigma^{1},\qquad d\sigma^{3}=0.$
A curl eigenfield will satisfy $\operatorname{curl}{v}=\alpha v$ for some $\alpha\in\mathbb{R}$, which is equivalent to writing
$dv^{\flat}=\alpha\star\!v^{\flat}.$
Now write $v$ in Fourier components, which must be (since the orbits of $\partial_{\theta}$ and $\partial_{z}$ both have length $2\pi$)
(4.19) $v^{\flat}=e^{in\theta}e^{imz}\left(if(r)\,\sigma^{1}+g(r)\,\sigma^{2}+h(r)\,\sigma^{3}\right),$
$dv^{\flat}$ $=e^{in\theta}e^{imz}\big(in\sigma^{2}+i(m-\tfrac{cn}{\varphi(r)^{2}})\sigma^{3}\big)\wedge\left(if(r)\,\sigma^{1}+g(r)\,\sigma^{2}+h(r)\,\sigma^{3}\right)$
$\qquad\qquad+e^{in\theta}e^{imz}\big(g^{\prime}(r)\,\sigma^{1}\wedge\sigma^{2}-h^{\prime}(r)\sigma^{3}\wedge\sigma^{1}\big)+e^{in\theta}e^{imz}\,\frac{2c\varphi^{\prime}(r)g(r)}{\varphi(r)^{3}}\,\sigma^{3}\wedge\sigma^{1}$
$=e^{in\theta}e^{imz}\bigg(\big(g^{\prime}(r)+nf(r)\big)\sigma^{1}\wedge\sigma^{2}+i\Big(nh(r)-(m-\tfrac{cn}{\varphi(r)^{2}})g(r)\Big)\sigma^{2}\wedge\sigma^{3}+$
$\qquad\qquad+\Big(-h^{\prime}(r)+\frac{2c\varphi^{\prime}(r)g(r)}{\varphi(r)^{3}}-\left(m-\frac{cn}{\varphi(r)^{2}}\right)f(r)\Big)\sigma^{3}\wedge\sigma^{1}\bigg)$
while the Hodge star formulas lead to
$\star v^{\flat}$ $=e^{in\theta}e^{imz}\left(i\varphi(r)\varphi^{\prime}(r)f(r)\,\sigma^{2}\wedge\sigma^{3}+\frac{\varphi^{\prime}(r)}{\varphi(r)}\,g(r)\,\sigma^{3}\wedge\sigma^{1}+\frac{\varphi(r)}{\varphi^{\prime}(r)}\,h(r)\,\sigma^{1}\wedge\sigma^{2}\right)$
Matching components then gives
$nh(r)-\left(m-\frac{cn}{\varphi(r)^{2}}\right)g(r)$ $=\alpha\varphi(r)\varphi^{\prime}(r)f(r)$
$g^{\prime}(r)+nf(r)$ $=\alpha\,\frac{\varphi(r)}{\varphi^{\prime}(r)}\,h(r)$
$h^{\prime}(r)-\frac{2c\varphi^{\prime}(r)g(r)}{\varphi(r)^{3}}+\left(m-\frac{cn}{\varphi(r)^{2}}\right)f(r)$ $=-\alpha\,\frac{\varphi^{\prime}(r)}{\varphi(r)}\,g(r).$
Using the first equation to eliminate $f$ (assuming $\alpha$ is nonzero), we get the system (4.15)–(4.16). Since the normal derivative at the boundary tori is $\partial_{r}$, the condition that $v$ is tangent to the boundary is equivalent to $f(a)=f(b)=0$, which translates into (4.17).
The only thing remaining is to lift $v^{\flat}$ given by (4.19) to a vector field. It is straightforward to compute from the metric that
The formula (4.14) then follows directly.∎
The simplest case is when $c=0$ and $\varphi(r)=r$ as in Remark 4.7, when the curl eigenfields reduce to Chandrasekhar-Kendall modes on the cylinder.
@@
When $c=0$, such as in the standard geometry on $S^{3}$ or the solid torus $D^{2}\times S^{1}$, we can get several explicit solutions when both $m$ and $n$ are nonzero.
In this case equations (4.15)–(4.16) become
$\alpha\varphi(r)\varphi^{\prime}(r)g^{\prime}(r)$ $=nmg(r)+\big(\alpha^{2}\varphi(r)^{2}-n^{2}\big)h(r)$
$\alpha\varphi(r)\varphi^{\prime}(r)h^{\prime}(r)$ $=\left(m^{2}-\alpha^{2}\varphi^{\prime}(r)^{2}\right)\,g(r)-nmh(r),$
with boundary condition nh(r)mg(r)=0nh(r)-mg(r)=0 at $r=a$ and $r=b$. If $\varphi(r)=s_{k}(r)$ so that $\operatorname{curl}{\partial_{z}}=2k\,\partial_{\theta}$ as in Remark 4.7, we can use the same technique as in Lemma 4.4 to get a five-dimensional space of vector fields preserved by curl, and compute eigenfields this way. However we will proceed in a more direct way for an even simpler case.
The simplest case is $k=0$ with $\varphi(r)=r$ on a solid torus with $a=0$ and $b=1$, where we can easily compute that $h$ satisfies the Bessel equation
$h^{\prime\prime}(r)+\frac{1}{r}\,h^{\prime}(r)+\Big(\alpha^{2}-m^{2}-\frac{n^{2}}{r^{2}}\Big)h(r)=0.$
@@
(4.20) $h(r)=-\beta^{2}J_{n}(\beta r),\qquad g(r)=\beta\alpha rJ_{n}^{\prime}(\beta r)+nmJ_{n}(\beta r),\qquad\alpha^{2}=\beta^{2}+m^{2},$
@@
$v(r,\theta,z)=e^{in\theta}e^{imz}\bigg(i\,\frac{nh(r)-mg(r)}{\alpha r}\,\partial_{r}+\frac{g(r)}{r^{2}}\,\partial_{\theta}+h(r)\,\partial_{z}\bigg).$
Hence the corresponding solution of the 3D Euler equation from Theorem 2.2, with $\lambda=2m/\alpha$, is given by
$u(t)=\partial_{\theta}+R\sin{\big(mz+n\theta+(\tfrac{2m}{\alpha}-n)\,t\big)}\big(m\beta J_{n}^{\prime}(\beta r)+\tfrac{n\alpha}{r}\,J_{n}(\beta r)\big)\,\partial_{r}\\ +R\cos{\big(mz+n\theta+(\tfrac{2m}{\alpha}-n)\,t\big)}\Big(\tfrac{1}{r^{2}}\big(\alpha\beta rJ_{n}^{\prime}(\beta r)+nmJ_{n}(\beta r)\big)\,\partial_{\theta}-\beta^{2}J_{n}(\beta r)\,\partial_{z}\Big),$
for any constant $R$. Kelvin originally derived these solutions in 1880 as solutions to the linearized Euler equation (i.e., valid only approximately for small $R$), and Dritschel [13] observed that they also solve the nonlinear Euler equation for any $R$.
On $S^{3}$ we can also use the same formulas from Example 4.5, where the curl eigenfields are already written in terms of the decomposition by $\partial_{\theta}$ and $\partial_{\psi}=\partial_{z}$, to obtain more examples of solutions to the $3D$ Euler equation.
@@
Choose $\varphi(r)=r$, a twisted version of the standard cylinder metric, with $c\neq 0$ and $0<a<b$. For $n=0$ and $m\in\mathbb{N}$, the system becomes
$-mg(r)$ $=\alpha rf(r)$
$g^{\prime}(r)$ $=\alpha rh(r)$
$h^{\prime}(r)-\frac{2cg(r)}{r^{3}}+mf(r)$ $=-\frac{\alpha g(r)}{r}$
with boundary conditions $g(a)=g(b)=0$. Eliminating $f$ and $h$, we get
$g^{\prime\prime}(r)=\frac{g^{\prime}(r)}{r}+\left(\frac{2c\alpha}{r^{2}}+m^{2}-\alpha^{2}\right)g(r),\qquad g(a)=g(b)=0.$
@@
$g(r)=c_{1}rJ_{\nu}(kr)+c_{2}rY_{\nu}(kr),\qquad k^{2}:=\alpha^{2}-m^{2},\qquad\nu^{2}:=2\alpha c+1,$
and there are nontrivial solutions vanishing at $r=a$ and $r=b$ if and only if
$J_{\nu}(ka)Y_{\nu}(kb)-J_{\nu}(kb)Y_{\nu}(ka)=0.$
@@
For example if $c=-\tfrac{3}{10}$ and $m=1$ with $a=\tfrac{2\pi}{3}$ and $b=2\pi$, then $\alpha=\tfrac{5}{4}$ works, since then $k=\tfrac{3}{4}$ and $\nu=\tfrac{1}{2}$, so that the Bessel functions become elementary. The corresponding solutions are multiples of
$f(r)=-4r^{-1/2}\cos{(\tfrac{3r}{4})},\qquad g(r)=5r^{1/2}\cos{(\tfrac{3r}{4})},\qquad h(r)=-3r^{-1/2}\sin{(\tfrac{3r}{4})}+2r^{-3/2}\cos{(\tfrac{3r}{4})}.$
Combining Theorem 4.9 and Theorem 2.2 to get $\lambda=\frac{2m}{\alpha}=\frac{8}{5}$, we conclude that for any $R$,
$u(t)=\partial_{\theta}+4Rr^{-1/2}\sin{\big(z+\tfrac{8}{5}t\big)}\cos{\big(\tfrac{3}{4}r\big)}\,\partial_{r}+5Rr^{-3/2}\cos{\big(z+\tfrac{8}{5}t\big)}\cos{\big(\tfrac{3}{4}r\big)}\,\partial_{\theta}\\ +R\cos{\big(z+\tfrac{8}{5}t\big)}\Big(-3r^{-1/2}\sin{\big(\tfrac{3}{4}r\big)}+2r^{-3/2}\cos{\big(\tfrac{3}{4}r\big)}\Big)\Big(\partial_{z}+\tfrac{3}{10r^{2}}\,\partial_{\theta}\Big)$
@@ 5. The Euler-Arnold Framework and Gerneralized Coriolis Force
The purpose of this section is to motivate Theorem 2.2 by exhibiting it as an application of a more general correspondence between Euler-Arnold dynamics and a class of modified equations incorporating a generalized Coriolis force.
@@
Note that $\operatorname{\operatorname{ad}^{\star}\!\!}$ is skew-adjoint with respect to $\langle\cdot,\cdot\rangle$ and therefore has purely imaginary spectrum on the complexified Lie algebra $\mathbb{C}\otimes\mathfrak{g}$. This observation will play a key role in the later construction of solutions.
@@
For ideal hydrodynamics, the condition $\operatorname{\operatorname{ad}^{\star}\!\!}_{X}=-\operatorname{ad}_{X}$ is equivalent to requiring that $X$ be a Killing field on the fluid domain. In this case, one may verify (at least locally in time) that the flow generated by $U$ is obtained by composing the flow of $V$ with the isometric flow generated by $X$, i.e. $\gamma_{U}=\gamma_{X}\circ\gamma_{V}$. On the round two-sphere, taking $X=\partial_{\theta}$, the term $\operatorname{\operatorname{ad}^{\star}\!\!}_{V}X$ reduces to the classical Coriolis force $2\Delta^{-1}\partial_{\theta}V$. This motivates our terminology of a generalized Coriolis force. Related ideas, including a notion of magnetic force, appear for geodesic motion on central extensions of diffeomorphism groups, cf. [2, 22, 35]. However here we work directly on the group and interpret the additional terms as arising from a change to a rotating frame generated by a Killing field.
@@
Surpressing $t$, we begin by taking a derivative in time of
@@
We now outline a mechanism for constructing solutions of (5.6), and hence solutions to the Euler–Arnold equation (5.2). The spectrum of the operator $z\mapsto\operatorname{\operatorname{ad}^{\star}\!\!}_{z}X$ naturally provides candidates for such solutions. Since it is skew-adjoint, its eigenvalues are purely imaginary and its eigenvectors generally lie in the complexified Lie algebra. This leads us to consider complex-valued solutions of (5.6). The following theorem shows that, under a simple algebraic condition, the real part of such a solution again yields a genuine real solution.
@@
Suppose that for some $T>0$ there exist $X\in\mathfrak{g}$ and a curve $Z\colon[0,T)\to\mathbb{C}\otimes\mathfrak{g}$ in the complexification of the Lie algebra solving (5.6). If $Z(t)=V(t)+iW(t)$ with $V(t),W(t)\in\mathfrak{g}$ such that
@@
$\partial_{t}Z=-\operatorname{\operatorname{ad}^{\star}\!\!}_{Z}Z-\operatorname{\operatorname{ad}^{\star}\!\!}_{Z}X$
@@
We now show how an eigenfield of the map $z\mapsto\operatorname{\operatorname{ad}^{\star}\!\!}_{z}X$ can be converted into a real solution of the Euler–Arnold equation. This result is the main algebraic ingredient in the explicit constructions presented in Sections 3 and 4.
@@
Let $X\in\mathfrak{g}$ with $\operatorname{\operatorname{ad}^{\star}\!\!}_{X}=-\operatorname{ad}_{X}$. If there exists $z\in\mathbb{C}\otimes\mathfrak{g}$ and $\lambda\in\mathbb{R}$ such that
$\operatorname{\operatorname{ad}^{\star}\!\!}_{z}X=-i\lambda z~,\quad\text{and}\quad\operatorname{\operatorname{ad}^{\star}\!\!}_{z}z=0,$
@@
(5.17) $Z(t)=\rho e^{i(\sigma+\lambda t)}z$
defines a complex-valued solution of (5.6). Moreover, if $z=v+iw$, with $v,w\in\mathfrak{g}$ and
@@
$\operatorname{\operatorname{ad}^{\star}\!\!}_{X}z=i\lambda z.$
@@
The fact that $Z(t)=\rho e^{i(\sigma+\lambda t)}z$ is a complex-valued solution of (5.6) is immediate. Writing $Z(t)=V(t)+iW(t)$, a direct computation gives
@@
The condition that $\operatorname{\operatorname{ad}^{\star}\!\!}_{z}z=0$ tells us that
@@
$\operatorname{\operatorname{ad}^{\star}\!\!}_{X}z=i\lambda z.$
@@
Hence, the condition $\operatorname{\operatorname{ad}^{\star}\!\!}_{z}X=-i\lambda z$ is equivalent to $K_{X}z=-i\lambda z$, while $\mathcal{A}z=\alpha z$ implies both $\operatorname{\operatorname{ad}^{\star}\!\!}_{z}z=0$ and $\operatorname{\operatorname{ad}^{\star}\!\!}_{w}w=0$. The added assumption of $X,z]=-\operatorname{ad}_{X}z=i\zeta z$ integrates to $\operatorname{Ad}_{\gamma_{X}(t)}z=e^{-i\zeta t}z$ and notably avoids the need for $X$ to be Killing, although all examples in Sections 3 and 4 fall into this category. Finally, the stationarity condition $\operatorname{\operatorname{ad}^{\star}\!\!}_{X}z=i\lambda z$ in Theorem 5.6 becomes $X,\mathcal{A}z]=i\lambda\mathcal{A}z$ which, under the inertia eigenvalue condition $\mathcal{A}z=\alpha z$, is equivalent to $X,z]=i\lambda z$. This recovers the condition $\lambda=\zeta$ appearing in Theorem 2.2.
@@
Accordingly, eigenfields $z$ with $\lambda=0$ produce solutions that are trivial in this moving-frame sense, whereas $\lambda\neq 0$ yields solutions $U$ that are genuinely non-trivial.
@@
The simplest nontrivial scenario occurs when the hierarchy truncates after the first-order term, i.e. when $x_{n}\equiv 0$ for all $n\geq 2$. In this case, the system reduces to

### v2

@@ 4.2.2. Curl eigenfields in a torus bundle
(4.11) $nb({\varepsilon_{1}})-\left(m-\frac{cn}{\varphi({\varepsilon_{1}})^{2}}\right)a({\varepsilon_{1}})=nb({\varepsilon_{2}})-\left(m-\frac{cn}{\varphi({\varepsilon_{2}})^{2}}\right)a({\varepsilon_{2}})=0,$
the constant $\alpha$ implicitly defined and $f$ given by
$f(r)=\,\frac{nb(r)-\big(m-\frac{cn}{\varphi(r)^{2}}\big)a(r)}{\alpha\varphi(r)\varphi^{\prime}(r)}.$
Then the complex fields
(4.12) $\mathfrak{q}(r,\theta,z)=e^{in\theta}e^{imz}\bigg(if(r)\,\partial_{r}+\frac{a(r)}{\varphi(r)^{2}}\,\partial_{\theta}+\frac{b(r)}{\varphi^{\prime}(r)^{2}}\Big(-\frac{c}{\varphi(r)^{2}}\,\partial_{\theta}+\partial_{z}\Big)\bigg)$
are curl-eigenfields satisfying the conditions of Theorem 2.2 with
$K_{\partial_{\theta}}\mathfrak{q}=-i\frac{2m}{\alpha}\mathfrak{q},\qquad\operatorname{curl}\mathfrak{q}=\alpha\mathfrak{q},\qquad[\partial_{\theta},\mathfrak{q}]=in\mathfrak{q}.$
In the simplest case, the curl eigenfields reduce to Chandrasekhar-Kendall modes on the cylinder.
@@
On the solid torus $D^{2}\times S^{1}$, with the standard flat metric $ds^{2}=dr^{2}+r^{2}\,d\theta^{2}+dz^{2}$ corresponding to $c=0$ and $\varphi(r)=r$ in (4.7), with ${\varepsilon_{1}}=0$ and ${\varepsilon_{2}}=1$, we can see that $b$ satisfies the Bessel equation
$b^{\prime\prime}(r)+\frac{1}{r}\,b^{\prime}(r)+\Big(\alpha^{2}-m^{2}-\frac{n^{2}}{r^{2}}\Big)b(r)=0.$
@@
(4.13) $b(r)=-\beta^{2}J_{n}(\beta r),\qquad a(r)=\beta\alpha rJ_{n}^{\prime}(\beta r)+nmJ_{n}(\beta r),\qquad\alpha^{2}=\beta^{2}+m^{2},$
@@
$v(r,\theta,z)=e^{in\theta}e^{imz}\bigg(i\,\frac{nb(r)-ma(r)}{\alpha r}\,\partial_{r}+\frac{a(r)}{r^{2}}\,\partial_{\theta}+b(r)\,\partial_{z}\bigg).$
Hence the corresponding solution of the 3D Euler equation from Theorem 2.2, with $\lambda=\frac{2m}{\alpha}$, is given by
$U(t)=\partial_{\theta}+\rho\sin{(mz+n\theta+\varpi t)}\big(m\beta J_{n}^{\prime}(\beta r)+\tfrac{n\alpha}{r}\,J_{n}(\beta r)\big)\,\partial_{r}\\ +\rho\cos{(mz+n\theta+\varpi t)}\Big(\tfrac{1}{r^{2}}\big(\alpha\beta rJ_{n}^{\prime}(\beta r)+nmJ_{n}(\beta r)\big)\,\partial_{\theta}-\beta^{2}J_{n}(\beta r)\,\partial_{z}\Big),$
for any constant $\rho\geq 0$, with $\varpi=\frac{2m}{\alpha}-n$. Kelvin originally derived these solutions in 1880 as solutions to the linearized Euler equation (i.e., valid only approximately for small $\rho$), and Dritschel [15] observed that they also solve the nonlinear Euler equation for any $\rho$.
@@
Choose $\varphi(r)=r$ in (4.7) with $c\neq 0$ and $0<{\varepsilon_{1}}<{\varepsilon_{2}}$. For $n=0$ and $m\in\mathbb{N}$, the system (4.9)–(4.11) becomes
$a^{\prime}(r)$ $=\alpha rb(r)$
$b^{\prime}(r)-\frac{2ca(r)}{r^{3}}+mf(r)$ $=-\frac{\alpha a(r)}{r}$
with boundary conditions $g({\varepsilon_{1}})=g({\varepsilon_{2}})=0$ and $\alpha rf(r)=-ma(r)$. Eliminating $b$, we get
$a^{\prime\prime}(r)=\frac{a^{\prime}(r)}{r}+\left(\frac{2c\alpha}{r^{2}}+m^{2}-\alpha^{2}\right)a(r),\qquad g({\varepsilon_{1}})=g({\varepsilon_{2}})=0.$
@@
$a(r)=c_{1}rJ_{\nu}(kr)+c_{2}rY_{\nu}(kr),\qquad k^{2}:=\alpha^{2}-m^{2},\qquad\nu^{2}:=2\alpha c+1,$
and there are nontrivial solutions vanishing at $r={\varepsilon_{1}}$ and $r={\varepsilon_{2}}$ if and only if
$J_{\nu}(k{\varepsilon_{1}})Y_{\nu}(k{\varepsilon_{2}})-J_{\nu}(k{\varepsilon_{2}})Y_{\nu}(k{\varepsilon_{1}})=0.$
@@
For example if $c=-\tfrac{3}{10}$ and $m=1$ with ${\varepsilon_{1}}=\tfrac{2\pi}{3}$ and ${\varepsilon_{2}}=2\pi$, then $\alpha=\tfrac{5}{4}$ works, since then $k=\tfrac{3}{4}$ and $\nu=\tfrac{1}{2}$, so that the Bessel functions become elementary. The corresponding solutions are multiples of
$f(r)=-4r^{-1/2}\cos{(\tfrac{3r}{4})},\qquad a(r)=5r^{1/2}\cos{(\tfrac{3r}{4})},\qquad b(r)=-3r^{-1/2}\sin{(\tfrac{3r}{4})}+2r^{-3/2}\cos{(\tfrac{3r}{4})}.$
Combining Theorem 4.12 and Theorem 2.2 to get $\lambda=\frac{2m}{\alpha}=\frac{8}{5}$, we conclude that for any $\rho\geq 0$,
$u(t)=\partial_{\theta}+4\rho r^{-1/2}\sin{\big(z+\tfrac{8}{5}t\big)}\cos{\big(\tfrac{3}{4}r\big)}\,\partial_{r}+5\rho r^{-3/2}\cos{\big(z+\tfrac{8}{5}t\big)}\cos{\big(\tfrac{3}{4}r\big)}\,\partial_{\theta}\\ +\rho\cos{\big(z+\tfrac{8}{5}t\big)}\Big(-3r^{-1/2}\sin{\big(\tfrac{3}{4}r\big)}+2r^{-3/2}\cos{\big(\tfrac{3}{4}r\big)}\Big)\Big(\partial_{z}+\tfrac{3}{10r^{2}}\,\partial_{\theta}\Big)$
@@ 5. The Euler-Arnold Framework and Gerneralized Coriolis Force
The purpose of this section is to motivate Theorem 2.2 and exhibit it as an application of a more general correspondence between Euler-Arnold dynamics and a class of modified equations incorporating a generalized Coriolis force.
@@
Note that for fixed $v\in\mathfrak{g}$ the operator $u\mapsto\operatorname{\operatorname{ad}^{\star}\!\!}_{u}v$ is skew-adjoint with respect to $\langle\cdot,\cdot\rangle$ and therefore has purely imaginary spectrum on the complexified Lie algebra $\mathbb{C}\otimes\mathfrak{g}$. This observation will play a key role in the later construction of solutions.
@@
The correspondence in Theorem 5.2 shows that the Euler-Arnold equation (5.2) and the modified equation (5.6) are equivalent up to a time-dependent change of variables. In particular, solutions of one system are in one-to-one correspondence with solutions of the other via (5.5). Thus, from a dynamical point of view, the two equations encode the same information: qualitative and quantitative properties of solutions (such as existence, regularity, or conserved quantities) may be transferred directly between them. In this sense, the additional term $\operatorname{\operatorname{ad}^{\star}\!\!}_{V}X$ can be interpreted not as introducing genuinely new dynamics, but rather as expressing the Euler-Arnold equation in a rotating frame generated by $X$.
Remark 5.4.
For ideal hydrodynamics, the condition $\operatorname{\operatorname{ad}^{\star}\!\!}_{X}=-\operatorname{ad}_{X}$ is equivalent to requiring that $X$ be a Killing field on the fluid domain. In this case, one may verify (at least locally in time) that the flow generated by $U$ is obtained by composing the flow of $V$ with the isometric flow generated by $X$, i.e., $\gamma_{U}=\gamma_{X}\circ\gamma_{V}$. On the round two-sphere, taking $X=\partial_{\theta}$, the term $\operatorname{\operatorname{ad}^{\star}\!\!}_{V}X$ reduces to the classical Coriolis force $2\Delta^{-1}\partial_{\theta}V$. This motivates our terminology of a generalized Coriolis force. Related ideas, including a notion of magnetic force, appear for geodesic motion on central extensions of diffeomorphism groups, cf. [2, 25, 40]. However here we work directly on the group and interpret the additional terms as arising from a change to a rotating frame generated by a Killing field.
@@
Suppressing $t$, we begin by taking a derivative in time of
@@
We now outline a mechanism for constructing solutions of (5.6), and hence solutions to the Euler–Arnold equation (5.2). The spectrum of the operator $\mathfrak{q}\mapsto\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathfrak{q}}X$ naturally provides candidates for such solutions. Since it is skew-adjoint, its eigenvalues are purely imaginary and its eigenvectors generally lie in the complexified Lie algebra. This leads us to consider complex-valued solutions of (5.6). The following theorem shows that, under a simple algebraic condition, the real part of such a solution again yields a genuine real solution.
@@
Suppose that for some $T>0$ there exist $X\in\mathfrak{g}$ and a curve $\mathcal{Q}\colon[0,T)\to\mathbb{C}\otimes\mathfrak{g}$ in the complexification of the Lie algebra solving (5.6). If $\mathcal{Q}(t)=V(t)+iW(t)$ with $V(t),W(t)\in\mathfrak{g}$ such that
@@
$\partial_{t}\mathcal{Q}=-\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathcal{Q}}\mathcal{Q}-\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathcal{Q}}X$
@@
We now show how an eigenfield of the map $\mathfrak{q}\mapsto\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathfrak{q}}X$ can be converted into a real solution of the Euler–Arnold equation. This result is the main algebraic ingredient in the explicit constructions presented in Sections 3 and 4.
@@
Let $X\in\mathfrak{g}$ with $\operatorname{\operatorname{ad}^{\star}\!\!}_{X}=-\operatorname{ad}_{X}$. If there exists ${\mathfrak{q}}\in\mathbb{C}\otimes\mathfrak{g}$ and $\lambda\in\mathbb{R}$ such that
$\operatorname{\operatorname{ad}^{\star}\!\!}_{{\mathfrak{q}}}X=-i\lambda{\mathfrak{q}}~,\quad\text{and}\quad\operatorname{\operatorname{ad}^{\star}\!\!}_{{\mathfrak{q}}}{\mathfrak{q}}=0,$
@@
(5.17) $\mathcal{Q}(t)=\rho e^{i(\sigma+\lambda t)}{\mathfrak{q}}$
defines a complex-valued solution of (5.6). Moreover, if ${\mathfrak{q}}=v+iw$, with $v,w\in\mathfrak{g}$ and
@@
$\operatorname{\operatorname{ad}^{\star}\!\!}_{X}{\mathfrak{q}}=i\lambda{\mathfrak{q}}.$
@@
The fact that $\mathcal{Q}(t)=\rho e^{i(\sigma+\lambda t)}{\mathfrak{q}}$ is a complex-valued solution of (5.6) is immediate. Writing $\mathcal{Q}(t)=V(t)+iW(t)$, a direct computation gives
@@
The condition that $\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathfrak{q}}{\mathfrak{q}}=0$ tells us that
@@
$\operatorname{\operatorname{ad}^{\star}\!\!}_{X}{\mathfrak{q}}=i\lambda{\mathfrak{q}}.$
@@
Hence, the condition $\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathfrak{q}}X=-i\lambda{\mathfrak{q}}$ is equivalent to $K_{X}{\mathfrak{q}}=-i\lambda{\mathfrak{q}}$, while $\mathcal{A}{\mathfrak{q}}=\alpha{\mathfrak{q}}$ implies both $\operatorname{\operatorname{ad}^{\star}\!\!}_{\mathfrak{q}}{\mathfrak{q}}=0$ and $\operatorname{\operatorname{ad}^{\star}\!\!}_{w}w=0$. The added assumption of $X,{\mathfrak{q}}]=-\operatorname{ad}_{X}{\mathfrak{q}}=i\zeta{\mathfrak{q}}$ integrates to $\operatorname{Ad}_{\gamma_{X}(t)}{\mathfrak{q}}=e^{-i\zeta t}{\mathfrak{q}}$ and notably avoids the need for $X$ to be Killing, although all examples in Sections 3 and 4 fall into this category. Finally, the stationarity condition $\operatorname{\operatorname{ad}^{\star}\!\!}_{X}{\mathfrak{q}}=i\lambda{\mathfrak{q}}$ in Theorem 5.7 becomes $X,\mathcal{A}{\mathfrak{q}}]=i\lambda\mathcal{A}{\mathfrak{q}}$ which, under the inertia eigenvalue condition $\mathcal{A}{\mathfrak{q}}=\alpha{\mathfrak{q}}$, is equivalent to $X,{\mathfrak{q}}]=i\lambda{\mathfrak{q}}$. This recovers the condition $\lambda=\zeta$ appearing in Theorem 2.2.
@@
Accordingly, eigenfields $\mathfrak{q}$ with $\lambda=0$ produce solutions that are trivial in this moving-frame sense, whereas $\lambda\neq 0$ yields solutions $U$ that are genuinely non-trivial.
@@
The simplest nontrivial scenario occurs when the hierarchy truncates after the first-order term, i.e., when $x_{n}\equiv 0$ for all $n\geq 2$. Remarkably, a sufficient condition for this is $\operatorname{\operatorname{ad}^{\star}\!\!}_{x_{1}}x_{1}=0$; indeed, then $x_{2}=0$ and all higher-order terms vanish inductively. In this case, the system reduces to